# A demonstration of the  class file for SAGE Publications

### v1

@@ 6.2 Task Planning Experiments
To evaluate the performance of the proposed task planning pipeline for clothes manipulation, we collect images annotated with ground-truth semantic keypoints (labeled by human experts) along with free-form natural language descriptions to evaluate the pipeline on folding, flattening, hanging, and placing tasks. For each task category, we collect 30 images of 15 diverse clothes. We evaluate the pipeline using two metrics: check completion accuracy, which assesses whether the pipeline correctly determines if the overall task has been completed based on the observation; and subtask accuracy, which measures whether the generated subtask sequence is appropriate for the observed state. To further analyze the role of semantic keypoints in task planning, we compare our pipeline (VLM + KP) with two baselines: (i) a VLM baseline that takes raw images without semantic keypoints, and (ii) an LLM baseline that receives only the language descriptors of semantic keypoints along with the language instruction, without any visual input. All pipelines receive the same few-shot demonstrations, but differ in visual input: our method uses keypoint-annotated images, the VLM baseline uses raw images, and the LLM baseline doesn’t rely on visual input.
@@
Table 2 presents the experimental results. Since the LLM baseline lacks visual input, we don’t report its check completion accuracy. It performs the worst on folding and flattening tasks, where adaptation to varied initial states is crucial. For the folding task, some parts of the clothes (such as sleeves) may already be folded. For flattening, the model needs to understand the clothes’ state from visual input and determine whether flattening is required. For example, in some flattening tasks, the robot must repeatedly flatten the clothes to achieve a flat state. The LLM baseline shows slightly better performance in hanging and placing tasks, which involve less uncertainty in the initial state. Moreover, the LLM’s symbolic reasoning remains consistent, robust, and unaffected by distribution shifts in visual input.
Our pipeline achieves reliable accuracy in proposing appropriate task plans and determining task completion. By providing semantic keypoints and a predefined skill library, we enable structured and controllable planning. To enhance robustness, we employ in-context learning with examples that include language instructions, observation images with semantic keypoint annotations, and corresponding task plans. All generated plans are also verified before execution.
@@
Compared to the VLM baseline without semantic keypoints, our pipeline achieves more reliable task planning, especially in long-horizon folding tasks. Without semantic keypoints, the VLM often hallucinates or generates wrong plans. Semantic keypoints provide critical structural information about the clothes’ state, especially under distribution shifts between observation images and the few-shot examples in the prompt. This allows the model to plan based on clothes’ state, rather than being misled by irrelevant visual differences.
@@
Table 3: Simulation Experiment Results. The average success rates (%) on testing tasks. The best performance is in bold.
@@ 6.3 Simulation Experiments
To evaluate CLASP’s performance on clothes manipulation tasks, we conduct experiments in SoftGym, where clothes are modeled as particles with ground truth positions. We sample 3D meshes of clothes from CLOTH3D (Bertiche et al., 2020) dataset, covering four common clothes types in human’s daily life: T-shirts, trousers, skirts, and towel. For each type, over 35 instances of varying sizes and shapes are used. In addition, we extend the SoftGym benchmark to 16 clothes tasks. These tasks can be divided into 4 types:
@@
We also implement an open-loop version of CLASP, where an LLM is prompted to generate a task plan based on the language instruction and semantic keypoint descriptors. In this version, task planning is performed only once at the beginning of the manipulation task and is not conditioned on the observation image.
@@
Figure 9: Real World Setup. We establish a dual-arm robot manipulation system for conducting real-world experiments on diverse clothes types.
Table 4: Real Experiment Results. The average success rates on various clothes and tasks.
@@ 6.4 Real-world Experiments
To evaluate the performance and generalization of CLASP in real-world scenarios, we establish a dual-arm robot manipulation system. As shown in Figure. 9, the system consists of two Franka Research 3 robot arms and a top-down RealSense L515 RGB-D camera for capturing RGB-D images. The clothes are placed on a table in front of the robot. We use SAM (Kirillov et al., 2023) and OWLv2 (Minderer et al., 2024) to segment the clothes for semantic keypoint extraction. To generate the dual-arm trajectories, we implement a motion planning algorithm using MoveIt! (Chitta et al., 2012) to avoid collisions and synchronize the motion between the arms.
@@
The experiment results are shown in Table 4. For certain clothes (#5, #13, #14, and #15), due to limitations in the robot’s working space, we did not perform evaluations on hanging and placing tasks. These are marked as N/A (Not Applicable). Overall, CLASP can be applied to a wide variety of clothes and manipulation tasks while achieving a high success rate. CLASP achieves an 86% success rate in clothes folding, a 66% success rate in clothes flattening, a 94% success rate in clothes hanging, and a 92% success rate in clothes placing. The success rate is comparable to existing task-specific clothes manipulation algorithms, while we test CLASP on broader clothes types and instances. With a general semantic keypoints representation and the prior knowledge from foundation models, CLASP is an effective general-purpose clothes manipulation method. Fig. 10 illustrates some representative examples. For each task, we visualize the extracted semantic keypoints, manipulation action sequences, and achieved object state. CLASP demonstrates the ability to manipulate various clothes in different ways in real-world scenarios. More videos of real-world experiments can be found in our supplementary material.
@@ 7 Discussion
Despite the promising experimental results, CLASP has several limitations: (1) Semantic keypoint extraction under significant occlusion: CLASP leverages existing vision foundation models and the VLM to extract semantic keypoints on open-category clothes without additional training. These models struggle to handle clothes with significant occlusion due to the limited representation of such data in their training datasets. Moreover, inferring the geometric structure of clothes under significant occlusion requires the model to possess a strong prior. This limitation contributes to the relatively lower success rate of CLASP in the clothes flattening task compared to other tasks. State estimation of deformable objects under significant occlusion remains an open challenge in robot perception. One potential improvement could involve designing heuristic policies to reduce the self-occlusion of clothes. This could involve identifying and properly separating different clothes layers or utilizing tools like Breeze to reduce occlusion (Xu et al., 2022). (2) Large clothes manipulation: CLASP struggles to generate executable subtasks for manipulating large clothes due to workspace constraints. This limitation could be addressed by integrating a mobile manipulator system (Yang et al., 2024) or using tools like a hook to pull distant clothes into the workspace (Toussaint et al., 2018). (3) Action execution: During the clothes manipulation process, sliding or deformation may influence the results. In CLASP, semantic keypoint extraction is based on large vision foundation models, making real-time semantic keypoint extraction impossible. One possible improvement could involve using distillation to train a smaller keypoint extraction model, which can track semantic keypoints during clothes manipulation and enable online trajectory optimization.
@@ Appendix A Discussion of Differences with the Conference Version
    Introduction of a new open-category semantic keypoint extraction pipeline, enhancing applicability and generalization (Section 4).
@@
A.1 Open-category Semantic Keypoint Extraction
In the conference version, we trained a model to extract semantic keypoints from observations. However, due to the high-dimensional configuration space of clothes, collecting sufficient data is challenging. To mitigate data scarcity, we imported 3D clothes meshes into the simulation for data augmentation. Nevertheless, the limited availability of 3D mesh models restricted the model to extracting semantic keypoints from only four types of clothes. In this manuscript, we propose a novel open-category semantic keypoint extraction pipeline (Sec. 4) that leverages large vision-language models and foundation vision models to overcome data scarcity and eliminate the need for training on labeled data. In this new pipeline, introducing a new clothes category requires only a prototype image—without any annotations—for autonomous semantic keypoint discovery (Sec. 4.1). This enhancement significantly improves the applicability and generalization of the proposed method.

### v2

@@ 6.2 Task Planning Experiments
To evaluate the performance of the proposed task planner for clothes manipulation, we collect images annotated with ground-truth semantic keypoints (labeled by human experts) along with free-form natural language descriptions to evaluate the planner on folding, flattening, hanging, and placing tasks. For each task category, we collect 30 images of 15 diverse clothes. We evaluate the planner using two metrics: check completion accuracy, which assesses whether the planner correctly determines if the overall task has been completed based on the observation; and subtask accuracy, which measures whether the generated subtask sequence is appropriate for the observed state. We compare our method (VLM + KP) with two baselines: (i) a VLM baseline that takes raw images without semantic keypoints, and (ii) an LLM baseline that receives only the language descriptors of semantic keypoints along with the language instruction, without any visual input (as in our previous work Deng and Hsu (2025)). All methods receive the same few-shot demonstrations but differ in visual input: our method uses keypoint-annotated images, the VLM baseline uses raw images, and the LLM baseline does not rely on any visual input.
@@
Table 2 presents the experimental results. Since the LLM baseline lacks visual input, we do not report its accuracy in checking completion. It performs the worst on folding and flattening tasks, where adaptation to varied states is crucial. For folding, some parts of the clothes (such as sleeves) may be folded or unfolded after execution. For flattening, the model needs to understand the state of the clothes from visual observation and determine whether flattening is required. For example, in some flattening tasks, the robot must repeatedly flatten the clothes to achieve a flat state. The LLM baseline, being an open-loop planner, generates task plans solely from symbolic descriptions without observing the outcomes of action execution. This open-loop nature leads to failures in tasks that require dynamic replanning. In contrast, our planner operates in a closed-loop manner, continuously updating the plan based on visual and keypoint feedback to adjust its strategy when action outcomes deviate from expectations. The LLM baseline performs slightly better in hanging and placing tasks, which involve less uncertainty in the initial state. Moreover, the LLM’s symbolic reasoning remains consistent and unaffected by visual distribution shifts.
@@
Compared to the VLM baseline without semantic keypoints, our pipeline achieves more reliable and interpretable task planning, especially in long-horizon folding tasks. Without semantic keypoints, the VLM often hallucinates or generates incorrect plans. Semantic keypoints provide critical structural information about the clothes’ state, allowing the model to reason over geometry and topology rather than surface appearance. This enhances robustness under distribution shifts between the observed images and the few-shot examples in the prompt.
Overall, our proposed planner achieves reliable accuracy in generating appropriate task plans and determining task completion. By providing semantic keypoints and a predefined skill library, we enable structured and controllable planning. To further enhance robustness, we employ in-context learning with examples that include language instructions, observation images with semantic keypoint annotations, and corresponding task plans. All generated plans are also verified before execution, and the planner supports dynamic replanning based on new observations, resulting in stable and robust closed-loop task planning.
@@
Table 3: Simulation experiments. The average success rates (%) on testing tasks. The best performance is in bold.
@@ 6.3 Simulation Experiments
To evaluate CLASP’s performance on clothes manipulation, we conduct experiments in SoftGym, where clothes are modeled as particles with ground-truth positions. We sample 3D meshes of clothes from CLOTH3D (Bertiche et al., 2020) dataset, covering four common clothes types in human’s daily life: T-shirts, trousers, skirts, and towel. For each type, over 35 instances of varying sizes and shapes are used. In addition, we extend the SoftGym benchmark to 16 clothes tasks. These tasks can be divided into 4 types:
@@
Figure 8: Robot experiments. The experimental setup consists of a Franka dual-arm system performing folding, flattening, hanging, and placing over many different clothes types.
Table 4: Robot experiments. The success rates on various clothes and tasks.
@@ 6.4 Real-world Experiments
To evaluate the performance and generalization of CLASP in real-world scenarios, we establish a dual-arm robot manipulation system. As shown in Figure 8, the system consists of two Franka Research 3 robot arms and a top-down RealSense L515 RGB-D camera for capturing RGB-D images. The clothes are placed on a table in front of the robot. We use SAM (Kirillov et al., 2023) and OWLv2 (Minderer et al., 2024) to segment the clothes for semantic keypoint extraction. To generate the dual-arm trajectories, we implement a motion planning algorithm using MoveIt! (Chitta et al., 2012) to avoid collisions and synchronize the motion between the arms.
@@
The experiment results are shown in Table 4. For clothes 5, 13, 14, and 15, due to limitations in the robot’s working space, we did not perform evaluations on hanging and placing tasks. These are marked as N/A (Not Applicable). Overall, CLASP can be applied to a wide variety of clothes and manipulation tasks while achieving a high success rate. CLASP achieves an 86% success rate in clothes folding, a 66% success rate in clothes flattening, a 94% success rate in clothes hanging, and a 92% success rate in clothes placing. The success rate is comparable to existing task-specific clothes manipulation algorithms, while we test CLASP on broader clothes types and instances. With a general semantic keypoints representation and the prior knowledge from foundation models, CLASP is an effective general-purpose clothes manipulation method. Fig. 9 illustrates some representative examples. For each task, we visualize the extracted semantic keypoints, manipulation action sequences, and achieved object state. CLASP demonstrates the ability to manipulate various clothes in different ways in real-world scenarios. More videos of real-world experiments can be found in our supplementary material.
@@ 7 Discussion
Despite the promising overall results, CLASP still shows several limitations based on the experimental results. First, CLASP struggles to extract accurate semantic keypoints of clothes with significant occlusion or deformation. This limitation arises because such cases are underrepresented in the datasets of the vision foundation models and VLMs that CLASP relies on. Although CLASP can perform flattening based on semantic keypoints to reduce occlusion, its success rate is lower than in other tasks (Table 3 and Table 4). Nevertheless, it achieves such performance and generalization without using any robot manipulation data. Future work could improve robustness by refining the policy through imitation learning with additional demonstrations. Second, CLASP faces challenges in manipulating large clothes due to workspace limitations. In real-world experiments, it sometimes fails to generate geometric feasible task plans for large clothes, preventing evaluation on certain hanging and placing tasks (e.g., clothes 5, 13, 14, and 15). This limitation could be addressed by integrating a mobile manipulator to expand the reachable workspace (Yang et al., 2024) or by adding dragging skills (Toussaint et al., 2018), allowing CLASP to handle larger clothes with random pose. Lastly, during manipulation, sliding or deformation can affect performance, especially for silk clothes (clothes 14 in real-world experiments). The folding success rate remains relatively low due to excessive sliding and deformation. This issue could be addressed by closed-loop action execution based on real-time semantic keypoint tracking. However, since CLASP’s keypoint extraction depends on large foundation models, real-time perception is still challenging. A practical improvement would be to use CLASP’s current pipeline to generate labeled data and distill a smaller, lightweight keypoint extractor, enabling real-time keypoint tracking and online trajectory optimization to enhance action robustness.
@@ Appendix A Discussion of Differences with the Conference Version
    Introduction of a new semantic keypoint extraction pipeline for open-category clothes, enhancing applicability and generalization (Section 4).
@@
A.1 Semantic Keypoint Extraction for Open-category Clothes
In the conference version, we trained a model to extract semantic keypoints from observations. However, due to the high-dimensional configuration space of clothes, collecting sufficient data is challenging. To mitigate data scarcity, we imported 3D clothes meshes into the simulation for data augmentation. Nevertheless, the limited availability of 3D mesh models restricted the model to extracting semantic keypoints from only four types of clothes. In this manuscript, we propose a novel semantic keypoint extraction pipeline for open-category clothes (Sec. 4). The proposed pipeline leverages large vision-language models and vision foundation models to overcome data scarcity and eliminate the need for training on labeled data. In this new pipeline, introducing a new clothes category requires only a prototype image—without any annotations—for autonomous semantic keypoint discovery (Sec. 4.1). This enhancement significantly improves the applicability and generalization of the proposed method.